OpenStudy (anonymous):

HELP PLEASE 12. (03.06 MC) Which equation can you write for a line that will pass through the points (3, 6) and (5, 18) in function notation. (5 points) f(x) = 2x + 12 f(x) = 6x − 12 y = 2x + 12 y = 6x − 12

OpenStudy (anonymous):

PLEASE HELP ME

OpenStudy (anonymous):

I will give u the answer to this one of u help me

OpenStudy (anonymous):

I WILL TRY TO HELP U BUT YOURS IT PRETTY LONG SO LET ME GIVE U TWO OF MY QUESTIONS DEAL I PROMISE

OpenStudy (anonymous):

Ok firs find the slope and tell me what u get

OpenStudy (anonymous):

Do u know how to find slope?

OpenStudy (anonymous):

YES I REPLIED TO UR POST

OpenStudy (anonymous):

NOW HELP ME

OpenStudy (anonymous):

Find the slope

OpenStudy (anonymous):

LOOK I AM REALLY TIRED AND REALLY DO NOT WANT TO DO WORK I MA SURE U CAN FIND SOMEONE ELSE AND SO CAN I TO HELP US SORRY BUT I WILL HELP U WITH 1 STE U PICK WHICH STEP

OpenStudy (anonymous):

I'm not asking u to help me anymore

OpenStudy (anonymous):

OK FINE I WAS JUST TRYING TO BE NICE BC U HELPED ME SORRY :(

OpenStudy (anonymous):

I just said that to find the answer to this problem u have to find the slope

OpenStudy (anonymous):

So did u get this problem already

OpenStudy (anonymous):

OK WELL I FORGET HOW TO FIND SLOPE WHAT ABOUT PART B

OpenStudy (anonymous):

Ok fine the answer is the second one or B

OpenStudy (anonymous):

And no it's ok I got it

OpenStudy (anonymous):

what i posted on your page about a problem i hope it helps u and bye

OpenStudy (anonymous):

Bye thnx

OpenStudy (anonymous):

hey i have two questions

OpenStudy (anonymous):

12. (03.06 MC) Which equation can you write for a line that will pass through the points (3, 6) and (5, 18) in function notation. (5 points) f(x) = 2x + 12 f(x) = 6x − 12 y = 2x + 12 y = 6x − 12 Score: 0 of 5 13. (03.07 LC) Choose the equation of the horizontal line that passes through the point (−5, 2). (5 points) y = −5 y = 2 x = −5 x = 2 Score: 0 of 5

OpenStudy (anonymous):

Ok for the fist one u find the slope

OpenStudy (anonymous):

and how i forget

OpenStudy (anonymous):

wait do u replace x with 0

OpenStudy (anonymous):

So.... 18-6/5-3= 6

OpenStudy (anonymous):

U do that^^

OpenStudy (anonymous):

what what u just posted or

OpenStudy (anonymous):

So the answer is the second one

OpenStudy (anonymous):

i meani kinda just gave u the answer to a big problem u had sooo

OpenStudy (anonymous):

Yes I told u the answer to the first question is the second one

OpenStudy (anonymous):

ok thanks but i need the steps written out to

OpenStudy (anonymous):

For the second one I think it is the third one x=-5

OpenStudy (anonymous):

again thx but need them worked out

OpenStudy (anonymous):

Ok the steps for the first one is..... 18-6/5-3 =12/2 =6 So the slope is 6 And then I know it's the se Lund one bc I had a problem just like that and I got it right so yeah hope that helped

OpenStudy (anonymous):

* second one

OpenStudy (anonymous):

And there r no steps for the second question I think.... Science it is asking for the horizontal

OpenStudy (anonymous):

and what is the correct answer again for the second one?

OpenStudy (anonymous):

It is either b or c I'm no 100% sure sorry but I think it is c

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

I hope I was able to help

OpenStudy (anonymous):

Np thank u

OpenStudy (anonymous):

How can I give u a medal I'm kinda new to this?